OpenStudy (anonymous):

In a school there are 200 students in year 11. 110 of these students are boys. What percentage of these students are girls explain how to work this out please

OpenStudy (anonymous):

Number of remaining students = 200 - 110 = 90 now 90 is what percent of 200 \[\frac{90}{200} \times 100 = 45 %\]

OpenStudy (anonymous):

thanks